An aquarium holding African cichlids is 2 inches higher than it is wide. Its length is 29 ​inches, and its volume is 10440 cubic inches.
What are the height and width of the​ aquarium?
Height= ___ inches.
Width ____ inches.

Answer by Boreal(15235)

width is x
height is x+2
length is 29
The volume is the product 29*x*(x+2)=29x^2+58x=10440
29x^2+58x-10440=0
divide by 29
x^2+2x-360=0
(x+20)(x-18)=0
x=18 inches width
height is 20 inches
That product is 360 sq inches
multiplied by 29 get 10,440 in^3.